OpenStudy (anonymous):

What is the vertex for f(x)=x^2 -4x-3?

OpenStudy (anonymous):

can't wait to see this...

OpenStudy (anonymous):

you want to complete the square. when you have an equation in the form: \[x^2+bx+c\]first you divide b by 2 and square the result: \[(-\frac{4}{2})^2 = 4\] You add and subtract that value to your function: \[x^2-4x+4-4-3\] Then you group and factor: \[(x^2-4x+4)-4-3 = (x-2)^2-7\] So the vertex is (-2, -7)

OpenStudy (anonymous):

...

OpenStudy (anonymous):

i knew it!

OpenStudy (anonymous):

It was 2,-7

OpenStudy (anonymous):

well(2,-7)

OpenStudy (anonymous):

whoops, sign error >.< its supposed to be what you said.

OpenStudy (anonymous):

no one want to use the oh so simple \[x=-\frac{b}{2a}=-\frac{-4}{2\times 1}=2\]

OpenStudy (anonymous):

(-b/2a,-D/4a) a = 1 ,b =-4 D = 16+12 = 28 => (2,-7)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i hate that formula!

OpenStudy (anonymous):

first coordinate is 2, second is what you get when you replace x by 2

OpenStudy (anonymous):

nothing could be simpler

OpenStudy (anonymous):

I prefer simple lol

OpenStudy (anonymous):

x = 2, y = 4-8-3=-7 done vertex is (2,-7) finito stick a fork in it

OpenStudy (anonymous):

>.>

OpenStudy (anonymous):

everyone here hates this simple simple simple formula

OpenStudy (anonymous):

no one objects to \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] but i you erase the radical part they have a fit. why is that/

OpenStudy (anonymous):

just get rid of the radical and you have \[-\frac{b}{2a}\] easy easy easy

OpenStudy (anonymous):

all that stupid completing the square work and keeping track of what you add and subtract? for the birds. it is always \[-\frac{b}{2a}\] and the explanation is simple too

OpenStudy (anonymous):

I cant stand when they throw all that extra crap in there, they're just going to have you take off

OpenStudy (anonymous):

exactly. the point is, if you know that the first coordinate is \[-\frac{b}{2a}\] then you know what the second coordinate is by substitution. you do not need to keep track all the way down the line